OpenStudy (anonymous):

HELP PLEASE. Let f(x) = 2x+3 and g(x) = x2 Part 1 [2 points] Find f(g(x)) Part 2 [2 points] Find g(f(x)) Part 3 [4 points] Use complete sentences to explain any difference in part 1 and part 2. I don't need the whole thing done, i just want to make sure i'm doing it right. For f(g(x)) i plugged in 2x(x^2) + 3. is that right? and how would i solve.

OpenStudy (jdoe0001):

$$ f(x) = 2x+3\\ g(x) = \color{red}{x^2}\\ f(g(x)) = 2(\color{red}{x^2})+3 $$

OpenStudy (jdoe0001):

there's really nothing to solve, or even simplify, is a simplified polynomial already

OpenStudy (anonymous):

I don't understand why it would be 2(x^2).. What happens to the x already with the 2? 2x(x^2) ?

OpenStudy (jdoe0001):

f(x) = 2x + 5(x +3) f( a ) = 2(a) + 5( (a) + 3) f( a + b) = 2(a + b) + 5( (a + b) + 3) f( g(x) ) = 2( g(x) ) + 5( (g(x)) + 3)

OpenStudy (jdoe0001):

so, the "x" in the HOST function, is really just a PLACEHOLDER for whatever is passed on as argument

OpenStudy (loser66):

got what he meant?

OpenStudy (anonymous):

I think so. At least, i understand what to plug in and solve, not that theres much solving.. So for the third part, it's asking for the difference between them, what's the difference? I dont really see any?

OpenStudy (anonymous):

i know they are kind of different in solving. The f(x) value replaces the x in g(x). But does that count as a difference? @jdoe0001 @Loser66

OpenStudy (jdoe0001):

yes

OpenStudy (loser66):

 you fill in the ? marks